OpenStudy (anonymous):

Plz help!!! I have a midterm on monday and have no idea how to solve this. I need help solving R2 and R3. We have to use only the concepts of circuits and can only use the formulas: V=IR and P=IV

OpenStudy (loser66):

have to see how the circuit is arranged. do you have a pic?

OpenStudy (anonymous):

oh sorry ya i thought i put the picture. just a sec

OpenStudy (anonymous):

if you cant see anything, just tell me and ill just write out the numbers

OpenStudy (loser66):

ok, so we need VIRP table, right?

OpenStudy (anonymous):

i guess

OpenStudy (loser66):

oh, Vincent is here, let him help. Please @Vincent-Lyon.Fr

OpenStudy (anonymous):

ok

OpenStudy (loser66):

for R2, you mark I =34A and V =40V , right?

OpenStudy (anonymous):

no just 40 V the 34A goes through the whole section since it has to be divided among the other resistors.

OpenStudy (anonymous):

right?

OpenStudy (loser66):

Please, post original picture with original given data. Don't put your stuff there.

OpenStudy (vincent-lyon.fr):

R2 = 40V / 34A = 1.18Ω etc.

OpenStudy (loser66):

Totally different!!! they ask you to calculate equivalent R between R2 and R3, right?

OpenStudy (loser66):

and where is R6? do you mix 2 problems ? hahahaha

OpenStudy (anonymous):

no. i just have to figure out the current and volts for each resistor and the battery with the given info. i had already figure everything out for all the resistors except R2 and R3

OpenStudy (anonymous):

that was for the next problem. sorry. ok here is the data for this problem

OpenStudy (anonymous):

sorry for the bad quality, im taking these with my ipod

OpenStudy (anonymous):

wait @loser66 where did you go? i still need help

OpenStudy (anonymous):

hey @Vincent-Lyon.Fr why did you divide 40V by 34A

OpenStudy (anonymous):

this is a parallel circuit right? the current has to distributed amongst the resistors so all the current cant go to one resistor right?

OpenStudy (vincent-lyon.fr):

R = V/I for a resistor.

OpenStudy (anonymous):

yes but not all the 34A goes to R2 since it has to go to R3 as well right?

OpenStudy (vincent-lyon.fr):

No, according to your picture 34 amps go through R2.

OpenStudy (anonymous):

then how much goes through R3?

OpenStudy (anonymous):

24? since 10 A would go to R4 and R5?

OpenStudy (anonymous):

plz help i have a midterm on monday!

OpenStudy (anonymous):

if you dont know how to solve it just say it so i can ask others

OpenStudy (anonymous):

dont just leave. thats not nice. thx for staying loser66

OpenStudy (loser66):

I am trying, but not have the answer yet. I need it, too. hihihi

OpenStudy (anonymous):

kk thx so much. i have the answers if you want but dont know how to solve it.

OpenStudy (loser66):

ok, just post. I need it, too.

OpenStudy (anonymous):

kk. hold on.

OpenStudy (loser66):

Thanks for posting, let me try to find the logic on it.

OpenStudy (anonymous):

no problem. the wierd thing is i have already solved it before with the right answers, but now i dont remembber how i did it. :( its so frustrating!!!1

OpenStudy (vincent-lyon.fr):

Since Vbat = 60 V, then V1 = 60 V too because they are in parallel. So I1 = 60/10 = 6A So I2 = 40 - 6 = 34A Since V3 = 20V, then V2 = V1 - V3 = 60 - 20 = 40 V So R2 = 40/34 = 1.17 ohm etc. use additivity of currents at nodes and additivity of potential differences in series. I must go.

OpenStudy (anonymous):

ok thx.